not sure if you can help...
I need to find dy/dx at the point (0,1), if y^5=(x+2)^4 + e^x Iny-15

Answer by amarjeeth123(570)

y^5=(x+2)^4 + e^x Iny-15
Differentiating on both sides with respect to x we get,
5y^4(dy/dx)=4(x+2)^3+e^x.lny+e^x/y*(dy/dx)
Substituting x=0 and y=1 in the above we get,
5(dy/dx)=4(2)^3+e^0.ln(1)+e^0/1*(dy/dx)
5(dy/dx)=32+1(0)+(dy/dx)
4(dy/dx)=32
(dy/dx)=8
Answer=8